12a-5b=18
15a-5b=30

Answer by nerdybill(7384)

12a-5b=18
15a-5b=30
.
Using the "elimination method", multiply the top equation by -1 and then combine both equations:
-12a+5b=-18
15a-5b= 30
-----------
3a = 12
a = 4
.
Using the above, substitute it back into either equation (I'll use the first one) and solve for 'b':
12a-5b=18
12(4)-5b=18
48-5b=18
-5b=-30
b= 6
.
Solution:
a = 4
b = 6